OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

I would use elimination, just because I am more comfortable solving questions that way. 2x + 5y = -3 -15x - 5y = - 10 2x*-15 + 5y*-15 = -3*-15 -15x*2 - 5y*2 = -10*2 -30x - 75y = 45 -30x - 10y = -20 Cancel -30x from both equations, -65y = 85 y = -85/65 Dividing by 5, y = - 17/13